Session: Nov 1998 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (D. Dallacasa at Fri Nov 20 18:58:44 MET 1998) Minor failures

Medicina (D. Dallacasa at Fri Nov 20 19:01:55 MET 1998) Minor failures
The observations went smoothly. There might have been a problem soon after the tape change. The FS was complaining about the head 2 positioning; this was fixed by re-entering the positioning by hand. A few seconds of the first forward pass on the second tape might have been lost.

Onsala (A. Polatidis at Sat Nov 21 11:44:03 MET 1998) Success
The problem with the recording head was fixed and we are observing again. No known problems.

Noto (C.Trigilio at Mon Nov 23 10:19:57 MET 1998) Success

Effelsberg (dgraham at Mon Nov 23 11:48:35 MET 1998) Success
seems OK

Westerbork (Tony Foley at Tue Nov 24 16:28:12 MET 1998) Minor failures
Possibly polarization wrong from 0800 to 0820UT. Many scans after 1130UT outside Hour Angle range. During a playback test tracks 26-32 gave high parity errors, but our decoder works badly at 8Mbits/s.

Jodrell (Alastair Gunn at Wed Nov 25 07:47:06 MET 1998) Minor failures
FS crash at start of scedule, okay by 08:07 UT. Tape didnt stop at 08:27 UT, error by PI in timing. Tape failed to start at 13:01 reversal, started manually at 13:05. Incorrect antenna.ctl file used, should be Mk2.

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:14:14 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: gb033a.antab. Read the comments in the file to get further information.